OpenStudy (anonymous):

Solve sqrt-99(sqrt-7 minus sqrt -13)

OpenStudy (anonymous):

\[\sqrt{-99}(\sqrt{-7} - \sqrt{-13})\] ?

OpenStudy (anonymous):

yes. @LogicalApple in terms of i. sorry.

OpenStudy (anonymous):

i was on my phone posting.

OpenStudy (anonymous):

wait, no @logicalapple \[\sqrt{-99}*(\sqrt{-7}-\sqrt{13})\]

OpenStudy (anonymous):

Oh I see. Try rewriting sqrt(-99) as i sqrt(99) and sqrt(-7) as i sqrt(7) Then distribute i sqrt(99): \[i \sqrt{99} (i \sqrt{7} - \sqrt{13} )\]

OpenStudy (anonymous):

\[3\sqrt{11}*(\sqrt{7}i - \sqrt{13})\] @logicalapple

OpenStudy (anonymous):

\[3i \sqrt{11} * (i \sqrt{7} - \sqrt{13}) = 3i^{2} \sqrt{77} - 3i \sqrt{143}\]